OpenStudy (anonymous):

I will fan and medal! What constant term should be used to complete the square? x^2 - 5x + _____ = 7 a) 49/25 b) 25/4 c) 4/25 d) -49/25

OpenStudy (anonymous):

u should checkout http://www.peeranswer.com , get help much faster :)

imqwerty (imqwerty):

x^2 - 5x +______ =7 x^2 - 5x +25/4 - 25/4 =7 (x-5/2)^2 =53/4 so to complete the square we need to add 25/4